OpenStudy (anonymous):

Find two consecutive positive integers such that the square of the second is added to 6 times the first is equal to 49????? HELP PLEASE

OpenStudy (anonymous):

trouble setting it up?

OpenStudy (anonymous):

what should the formula look like?

OpenStudy (anonymous):

well, there are two numbers to look for... so two variables... which means you need two equations to solve it

OpenStudy (anonymous):

let X be the first number and Y be the second number, can you dig an equation out of that sentence?

OpenStudy (anonymous):

X and Y are consecutive integers. What does that mean in terms of X and Y? X = what?

Directrix (directrix):

Let x be a positive integer. Then, x +1 is the next consecutive positive integer 6x + (x+1)^2 = 49 6x + x^2 + 2x + 1 = 49 x^2 +8x -48 =0 (x +12)(x- 4)=0 Zero Product Property x = -12 and x = 4 Integers must be positive so -12 is not considered. Answer: 4 and 5

OpenStudy (anonymous):

thank you